### v2

# Scalp Diagnostic System With Label-Free Segmentation and Training-Free Image Translation

Abstract
Scalp diseases and alopecia affect millions of people around the world, underscoring the urgent need for early diagnosis and management of the disease. However, the development of a comprehensive AI-based diagnosis system encompassing these conditions remains an underexplored domain due to the challenges associated with data imbalance and the costly nature of labeling. To address these issues, we propose “ScalpVision”, an AI-driven system for the holistic diagnosis of scalp diseases and alopecia.
In ScalpVision, effective hair segmentation is achieved using pseudo image-label pairs and an innovative prompting method in the absence of traditional hair masking labels. This approach is crucial for extracting key features such as hair thickness and count, which are then used to assess alopecia severity. Additionally, ScalpVision introduces DiffuseIT-M, a generative model adept at dataset augmentation while maintaining hair information, facilitating improved predictions of scalp disease severity. Our experimental results affirm ScalpVision’s efficiency in diagnosing a variety of scalp conditions and alopecia, showcasing its potential as a valuable tool in dermatological care.
1 Introduction
Approximately 80 million people in the United States are affected by male and female pattern hair loss [36]. Additionally, a study by Elewski et al. [14] reveals that nearly $90\%$ of adults in the U.S. experience various scalp disorders. Among these, alopecia is especially common, impacting a significant portion of the population. The causes of alopecia are diverse, including aging, genetics, stress, and overall health conditions [46]. Therefore, early diagnosis is key to preventing the progression of these scalp-related diseases [35, 37].
Recognizing the importance of early detection, numerous studies have explored scalp disease diagnosis using microscopic scalp imagery. For example, Chang et al. [7] proposed a scalp diagnostic system, Kim et al. [24] concentrated on follicle detection, and Seo et al. [43] identified features of hair loss for diagnosing alopecia. However, these studies often treat scalp and hair issues separately, overlooking a more integrated diagnostic approach.
Figure 1: Data distribution and examples of AI-Hub microscopic scalp imaging training dataset. The images on the left side represent examples of each scalp disease condition at different severity and the charts on the right side represent data distribution of different severity within each scalp condition.
Figure 2: ScalpVision pipeline overview. Here, $I$ denotes the original image. The hair segmentation mask generated by the segmentation model $S$, which is trained using a pseudo-training set, is represented as $\hat{M}$. The mask produced by the SAM is denoted as $M_{\text{AP}}$, and $M$ signifies the final, combined hair segmentation mask. The “Automatic Prompt”, used for refining segmentation, is derived from $\hat{M}$.
Acquiring comprehensive labels for medical diagnostic systems presents a significant challenge in medical research. This difficulty primarily arises from the need for specialized domain knowledge in image annotation, which significantly increases labeling costs. Furthermore, data imbalance is a frequent issue, largely due to biases in the training process. For instance, most datasets focusing on scalp conditions disproportionately represent healthy scalps, leading to a scarcity of data on severe scalp diseases and alopecia. As a result, there are no publicly accessible datasets that offer detailed labels for both scalp conditions and alopecia in scalp images.
Despite the lack of publicly available datasets, we accessed a specialized dataset from AI-hub [1] for classifying the severity of scalp conditions and alopecia  [^1] This dataset is provided by ‘The Open AI Dataset Project (AI-Hub, S. Korea)’ and is exempt from IRB approval as it does not contain any information that can identify individuals.. This dataset categorizes the severity of scalp conditions, including dandruff, excess sebum, and erythema, into four levels (good, mild, moderate, severe) across 96K images. It also includes about 24K images for alopecia severity classification. The distribution of this dataset’s attributes is illustrated in Figure 1. However, this dataset lacks detailed information on hair and follicles, which is crucial for more in-depth diagnostic analyses.
To address these challenges, we developed “ScalpVision,” a comprehensive system designed for in-depth assessment of scalp health, as illustrated in Figure 2. ScalpVision’s pipeline integrates three key components: 1) hair segmentation, 2) diagnosis of scalp conditions, and 3) prediction of alopecia severity. The backbone of ScalpVision is its advanced segmentation technique. This technique employs label-free learning and zero-shot segmentation to effectively discern and extract hair-related features. It combines masks generated by a model trained on synthetic image-label pairs with those from the point-guided Segment Anything Model (SAM) [26]. For diagnosing scalp conditions, we leverage DiffuseIT-M, a diffusion-based image-to-image translation model, adept at addressing data imbalance. Finally, in the alopecia severity assessment phase, we utilize hair masks to derive features like hair thickness and count, which are then analyzed using machine learning models to provide accurate evaluations.
In summary, our main contributions are as follows:
  • We present a comprehensive diagnostic pipeline for microscopic scalp imagery that assesses both scalp condition and alopecia severity.
  • We develop a novel segmentation approach, combining heuristic-driven pseudo-labels with SAM to effectively segment images in the absence of ground truth labels.
  • We introduce DiffuseIT-M, a mask-guided, diffusion-based image-to-image translation method, facilitating targeted control over specific image regions without additional training.
We contribute to the field of scalp diagnosis by sharing our code on GitHub  [^2] https://github.com/winston1214/ScalpVision.
2 Related Works
Scalp image analysis. In the field of scalp image analysis, hair is a pivotal feature, often utilized for segmentation purposes [44, 48]. Investigations into the number [52, 41] and thickness [23] of hairs on microscopic scalp images have been conducted. Moreover, various approaches utilize microscopic scalp images to diagnose diseases. Systems for classifying scalp diseases and their severity have been developed [7, 25], while others assess alopecia severity using scalp information [24, 43].
Image augmentation. Previous studies have proposed methods for data augmentation by merging images [53, 54] or losing partial information of images [58, 8]. Recent developments in generative models, particularly the diffusion models [20, 40], have led to the proposal of image augmentation methods using synthetic images. Azizi and Trabucco et al. [3, 47] show diffusion models can generate photo-realistic images with given conditioning variables and improve image classification accuracy, even in medical applications [11, 57].
Label-free image segmentation on medical image. The challenges and expenses associated with acquiring medical image data have spurred the development of data-efficient training methods [9] and label-free segmentation models. Kim et al. [22] introduced a synthetic angiogram based on fractal design as a label, training a diffusion module for vessel segmentation. Large pre-trained segmentation models like the Segment Anything Model (SAM) [26] have garnered interest and demonstrated promise across various vision tasks [59]. While these models are increasingly applied to specific medical imaging challenges, success has been limited [17].
3 Our Approach: ScalpVision
To address the challenges of limited and imbalanced datasets in scalp imaging, we introduce ScalpVision, a diagnostic system designed for accurately assessing the severity of scalp-related diseases and alopecia. Central to ScalpVision is a hair segmentation module (Section 3.1), pivotal for both alopecia severity prediction (Section 3.2) and generating diverse scalp images to augment training datasets for scalp condition classification (Section 3.3). The overall pipeline of ScalpVision is illustrated in Figure 2.
3.1 Label-Free Hair Segmentation
For the precise diagnosis of alopecia and scalp conditions, our initial step involves segmenting hair within microscopic scalp images. However, as the dataset does not provide explicit segmentation labels, conventional supervised learning methods are not feasible.
Heuristic-driven pseudo-labeling. To address this, we first generate pseudo labels for training our segmentation model (S as shown in Figure 2) using prior knowledge. With the intuition that the hair on the microscopic scalp images follows either a linear function or a power function, we generate synthetic images to effectively guide the model to learn hair patterns on the scalp images. For each disease condition, we randomly select one image representing each distinct severity level, extract three smaller patches from regions of the scalp with no visible hair, and draw curves to simulate hair patterns. Additionally, to simulate dandruff noise, circular white shapes are added to these patches, but are not indicated in the pseudo masks, thus training the model to interpret them as noise. We generate $3{,}000$ pseudo-images and corresponding pseudo mask labels, using them to train the U2-Net [39]. This training generates the binary mask, represented as:
$\hat{M}=[\hat{M}(i,j)]\in\{0,1\}^{H\times W},$ (1)
where $H$ and $W$ are the height and width of the image, and $i\in[1,H]$, $j\in[1,W]$ denote pixel coordinates.
Algorithm 1 Extraction of representative points from mask
Input: Mask $\hat{M}$, bounding box size $n$, and cross-shaped structuring element kernel
Output: Representative hair points from mask $\hat{\mathcal{C}}$
1:  $\mathcal{H}_{\text{copy}}$leftarrow $\hat{{M}}$
2:  ^$\hat{\mathcal{H}}_{\text{skel}}$leftarrow zero array with same size as $\mathcal{H}_{\text{copy}}$
3:  while $\mathcal{H}_{\text{copy}}$$\neq$00 do
4:     Eroded, Dilated $\leftarrow$ MORPHOLOGY(((($\mathcal{H}_{\text{copy}}$, kernel))))
5:     $\hat{K}$leftarrow $\mathcal{H}_{\text{copy}}$- Dilated
6:     ^$\hat{\mathcal{H}}_{\text{skel}}$leftarrow ^$\hat{\mathcal{H}}_{\text{skel}}$lor $\hat{K}$
7:     $\mathcal{H}_{\text{copy}}$leftarrow Eroded
8:  end while
[…]
10:  for all  $(x,y)$in $\mathcal{H}_{\text{skel}}$ do
11:     $\hat{B}$leftarrow $\hat{B}$cup $\{(x-\frac{1}{2}n,y-\frac{1}{2}n,x+\frac{1}{2}n,y+\frac{1}{2}n)\}$
12:  end for
13:  $\hat{B}$leftarrow NMS($\hat{B}$)
[…]
15:  for all $(x_{1},y_{1},x_{2},y_{2})$in $\hat{B}$ do
[…]
17:  end for
18:  return $\hat{\mathcal{C}}$
Automatic prompting for SAM. To refine the hair segmentation mask $\hat{M}$, we utilize the foundation segmentation model, SAM [26], employing a point-prompting method to differentiate hair from scalp without additional training. However, we observed that selecting random points from $\hat{M}$ for positive point prompts often resulted in suboptimal masks. This issue was primarily due to points near the edges of $\hat{M}$ causing confusion for the SAM. Moreover, due to inherent randomness, there were instances in which sampled points clustered in a small localized region, resulting in the SAM segmenting only a limited number of hairs. To address these issues, we have developed an automatic prompting method. This method ensures uniform sampling across $\hat{M}$ and provides high-confidence guidance to the SAM based on the coarse segmentation mask $\hat{M}$. The specific steps of this method are detailed in Algorithm 1.
To extract the distinct features of the hair, we compute the skeletonized mask ^$\hat{\mathcal{H}}_{\text{skel}}$ ($\in\{0,1\}^{H\times W}$) using morphological erosion and dilation following [56]. This method operates by iteratively processing the image, eroding and dilating pixels at the object’s edges until no further pixel removal is possible. Then, we generate bounding boxes around each pixel in ^$\hat{\mathcal{H}}_{\text{skel}}$ with size $n\times n$ where we set $n=10$. These boxes undergo non-maximum suppression (NMS) to filter out the bounding boxes, denoted as $\hat{B}=\{\hat{b}_{j}\}_{j=1}^{k}$, where each box is defined by coordinates $(x_{\text{min}},y_{\text{min}},x_{\text{max}},y_{\text{max}})$. Following this, the mean points of the hair pixels, $\hat{\mathcal{C}}=\{\hat{c}_{j}\}_{j=1}^{k}$, in each bounding box $\hat{B}$ can be determined. For each $\hat{b}_{j}=(x_{1},y_{1},x_{2},y_{2})$, the mean point $\hat{c}_{j}=(\bar{x},\bar{y})$ is given by:
$\bar{x}=\frac{\sum_{i}\sum_{j}i\cdot\hat{\mathcal{H}}(i,j)}{\sum_{i}\sum_{j}\hat{\mathcal{H}}(i,j)},\quad\bar{y}=\frac{\sum_{i}\sum_{j}j\cdot\hat{\mathcal{H}}(i,j)}{\sum_{i}\sum_{j}\hat{\mathcal{H}}(i,j)}$ (2)
where the summation is over all $i\in[x_{1},x_{2}]$ and $j\in[y_{1},y_{2}]$.
Subsequently, we select positive point prompts for SAM from the calculated mean points $\hat{\mathcal{C}}$. For the negative point prompts, we utilize the inverse of the initial mask, specifically $1-\hat{M}$. These prompts, automatically generated, guide SAM in generating the binary segmentation mask $M_{\text{AP}}$ ($\in\{0,1\}^{H\times W}$).
Mask ensemble. $M_{\text{AP}}$ and $\hat{M}$ complement each other with strengths and weaknesses. $\hat{M}$ is robust against noise like dandruff as it was trained using simulated noise. Meanwhile, $M_{\text{AP}}$, benefiting from SAM’s superior edge detection, excels in constructing a clear boundary between hair and scalp. Therefore, to make a robust hair mask, the final binary mask, $M$, is derived from $\hat{M}$ and $M_{\text{AP}}$ with the AND operation, as follows:
$M=\hat{M}\land M_{\text{AP}}.$ (3)
3.2 Alopecia Severity Prediction
Alopecia primarily manifests through two indicators: hair thinning and follicle reduction [38]. Consequently, hair thickness and count are crucial metrics for assessing its severity. However, lacking ground-truth labels for these features, we have developed straightforward yet effective methods to approximate hair thickness (Figure 3(a)) and hair count (Figure 3(b)). These methods utilize the hair mask $M$ generated earlier, from which we extract bounding boxes $B$, and their respective mean points $C$, following the approach described in Section 3.1.
Figure 3: Feature extraction for alopecia severity prediction.
Hair thickness. Our objective is to measure the thickness of hair within each bounding box in the set $B=\{b_{j}\}_{j=1}^{k}$. Firstly, we define the set of pixel coordinates for hair within a bounding box $b_{j}$ as:
$\mathcal{P}_{j}=\{(x,y)\in b_{j}\mid M(x,y)=1\},$ (4)
where $M(x,y)=1$ indicates the presence of hair at the pixel coordinates $(x,y)$.
Next, we calculate $\mathcal{S}=\{s_{j}\}_{j=1}^{k}$ where each $s_{j}$ represents the slope perpendicular to the hair in the $j$-th bounding box. This slope is derived from the second principal component in a Principal Component Analysis (PCA) of the pixel coordinates $\mathcal{P}_{j}$. A perpendicular line ($\mathcal{L}_{j}$) is then defined using this slope $s_{j}$ and the mean point $c_{j}$:
$\mathcal{L}_{j}=\{(x,y)\mid s_{j}(x-c_{j,x})+c_{j,y}=y\}.$ (5)
Following this, we identify two boundary points of the hair, $p_{j,1}$ and $p_{j,2}$, located where $\mathcal{L}_{j}$ intersects with the edge points of $\mathcal{P}_{j}$. Denoting these edge points as $E(\mathcal{P}_{j})$, the boundary points are formulated as:
$p_{j,1},p_{j,2}=E(\mathcal{P}_{j})\cap\mathcal{L}_{j}.$ (6)
Finally, the pixel-level hair thickness ($t_{j}$) is calculated as the Euclidean distance between these two boundary points:
$t_{j}=||p_{j,1}-p_{j,2}||_{2}.$ (7)
Thus, we obtain a set of hair thicknesses: $\mathcal{T}=\{t_{j}\}_{j=1}^{k}$.
Hair count. To quantify the number of hairs, we utilize the Hough transformation [13], an established technique for line detection through parameter space representation, applied to the hair skeleton $\mathcal{H}_{\text{skel}}$. Inspired by the approach in [45], we enhance the method’s effectiveness for curved hair by implementing the Hough transformation at three different resolutions: original, half, and double. At each of these resolutions, the transformation identifies a set of line segments, where each segment is defined by a pair of endpoints, denoted as $\{(x_{1},y_{1}),(x_{2},y_{2})\}$. The line segments obtained from each resolution level are then resized to correspond to the original resolution scale and collectively merged. During this merging phase, we identify and remove duplicate line segments. This elimination is based on their similarity in vector angle, length, and spatial positioning.
The two features we have predicted, hair thickness and count, not only hold potential for direct use by clinicians in diagnosing alopecia but also present an opportunity for further exploration. Considering the dataset includes labels on the severity of alopecia, we investigate the application of these features as inputs for traditional machine learning models. This approach is aimed at enhancing the prediction of alopecia severity using the available dataset.
Figure 4: Overview of DiffuseIT-M. $x_{src}$ and $x_{trg}$ represent the source image and target image, respectively, with $M$ denoting the mask image of $x_{src}$. In the reverse process, we guide the generation to adhere to the conditions using the objective function. $\odot$ signifies the weighted sum of two images. “DINO” means DINO-ViT.
3.3 Scalp Condition Classification
Accurately classifying the severity of scalp-related diseases in microscopic images is challenging due to the rarity of extreme cases, as illustrated in Figure 1. To address this issue, we have developed a method that translates given images into scalp images of a desired condition. This method, named DiffuseIT-M, is a diffusion-based image translation model with mask guidance. It is designed to preserve the content (e.g., hair) of the source image while enabling the transformation into various scalp conditions. DiffuseIT-M is based on DiffuseIT [27] and incorporates an image editing technique inspired by Blended diffusion [2] to ensure hair preservation. The overview of DiffuseIT-M is depicted in Figure 4.
Preliminary. In the forward diffusion process, we adopted the DDPM [20] forward process. Given a clean image, $x_{0}\sim q(x_{0})$, Gaussian noise is gradually added at every time steps, $t$, using the Markov chain.
$q(x_{1:T}|x_{0}):=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),$ (8)
following the formulations and notations in [33]. This equation can be simplified in terms of $x_{t}$ sampled in one step as follows:
$x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (9)
where $\epsilon\sim\mathcal{N}(0,\bf{I})$, $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$.
In the reverse diffusion process, $q(x_{t-1}|x_{t})$ is approximated as $p_{\theta}(x_{t-1}|x_{t})$ with the trainable parameter $\theta$. Therefore, it is modeled by a Markov chain with a learned mean and fixed variance, initiated from $p(x_{T})$:
$p_{\theta}(x_{0:T}):=p_{\theta}(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t})$ (10)
where $p_{\theta}(x_{t-1}|x_{t}):=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. Rather than inferring $\mu_{\theta}(x_{t},t)$ directly, it predicts the noise $\epsilon_{\theta}(x_{t},t)$ added to $x_{0}$ to obtain $x_{t}$. Then, ${\mu}_{\theta}(x_{t},t)$ is derived using Bayes’ theorem:
${\mu}_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\Big{)}.$ (11)
Then, the sampling from the reverse diffusion is defined as:
$x_{t-1}={\mu}_{\theta}(x_{t},t)+\sigma_{t}\epsilon.$ (12)
Image translation with mask guidance. To facilitate the transfer of scalp disease characteristics while preserving hair features in our model, we utilize a comprehensive loss function, $\ell_{total}$. This function guides the reverse process and is composed of five distinct loss components. These components consider the source image ($x_{src}$), the target image ($x_{trg}$), and the hair mask ($M$) as inputs. The combined loss function is defined as:
$\begin{gathered}\ell_{total}(x;x_{src},x_{trg},M)=\lambda_{1}\ell_{style}+\lambda_{2}\ell_{content}\\ +\lambda_{3}\ell_{mask}+\lambda_{4}\ell_{sem}+\lambda_{5}\ell_{rng},\end{gathered}$ (13)
where $\lambda_{i\in[1,5]}$ denotes the weights assigned to each of these loss functions.
For $\ell_{style}$ and $\ell_{content}$, we utilize the style and content loss functions from DiffuseIT. We employ the [CLS] token matching loss using DINO ViT [6] to reflect semantic information in $x_{trg}$ and use keys of multi-head self-attention layers to preserve the content of $x_{src}$.
Additionally, to ensure hair preservation while translating scalp styles, we construct mask preservation loss function as:
$\ell_{mask}=\mbox{\small{LPIPS}}(x_{src}\odot M,\hat{x}_{0}(x_{t})\odot M)+||(x_{src}-\hat{x}_{0}(x_{t}))\odot M||_{2},$ (14)
where LPIPS denotes the learned perceptual image patch similarity metric [55].
We also include two other loss functions: $\ell_{rng}$, representing the squared spherical distance as proposed in [10], and $\ell_{sem}$, indicating the semantic divergence loss as outlined in [27].
Using this composite loss function, $\ell_{total}$, we guide the generation of the next sample step, $x_{t-1}$. To maintain hair information, we apply masking to the images:
$x_{t-1}=x_{t}\odot(1-M)+\Big{(}\hat{x}_{0}(x_{t})-\nabla_{x_{t}}\ell_{total}(\hat{x}_{0}(x_{t}))\Big{)}\odot M,$ (15)
where $\hat{x}_{0}(x_{t})$ is the estimation of the cleaned image derived from the sample $x_{t}$. This estimation is computed as follows:
$\hat{x}_{0}(x_{t})=\frac{x_{t}}{\sqrt{\bar{\alpha}_{t}}}-\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$ (16)
With this guide, we can translate the image style of the scalp without the need for additional training.
Classification strategy. Utilizing our DiffuseIT-M, as described above, we enhance our training set by translating randomly chosen images into ones with higher severity levels. We used random selection for the source images and weighted sampling for target images, where the likelihood of choosing an image was inversely proportional to the number of samples in its severity class.
For our classification task, we employed a pretrained backbone model and fine-tuned it with four Multi-Layer Perceptron (MLP) heads. Each head is associated with a specific loss function. The first head is designed to identify the presence of each scalp disease (i.e., dandruff, excess sebum, and erythema) without considering their severity. The remaining three heads are tasked with classifying the severity levels of each disease, categorized into four stages: good, mild, moderate, and severe. Consequently, our overall classification objective, denoted as $\ell_{cls}$, is expressed by the following formula:
$\ell_{cls}=\ell_{dis}+\ell_{dand}+\ell_{seb}+\ell_{ery},$ (17)
where $\ell_{dis}$ is the binary cross-entropy loss for the disease presence classification head, and $\ell_{dand}$, $\ell_{seb}$, and $\ell_{ery}$ represent the cross-entropy losses associated with determining the severity levels of each disease.
Figure 5: Comparison of various segmentation methods on hair. “GT” represents the mask images for which we have manually annotated the pixel segmentation. Note that $\hat{M}$, $M_{\text{AP}}$ and $M$ are proposed in Section 3.1.
4 Experiments
To showcase the effectiveness of our pipeline, we conducted a series of quantitative and qualitative experiments for each component of the pipeline.
4.1 Dataset
We trained and evaluated our model using the scalp image dataset from AI-Hub [1], previously mentioned in Section 1 and depicted in Figure 1. The dataset comprises $95{,}910$ images with a resolution of $640\times 480$ pixels from $20{,}000$ patients. It is divided into $72{,}342$ images for training and $23{,}568$ images for testing, with a subset of $21{,}703$ images from the training set designated for validation. Dermatologists classified these scalp images into one or more of the following conditions: dandruff, excess sebum, and erythema. Each condition was further assessed for severity, categorized as mild, moderate, or severe, with an additional category, good, indicating no presence of the condition. Furthermore, a subset of $23{,}803$ images, divided into $12{,}902$ for training, $5{,}613$ for validation, and $5{,}288$ for testing, includes specific labels for alopecia severity, following the same four-level categorization.
4.2 Hair Segmentation
Table 1: Performance of Hair segmentation on manually annotated test set.
Approach Pixel-F1 Jaccard Dice
Shih et al. [44] $0.7063$ $0.3475$ $0.5119$
Yue et al. [52] $0.7943$ $0.4933$ $0.6540$
Kim et al. [23] $0.8150$ $0.5614$ $0.7078$
SAM [26] $0.5030$ $0.3609$ $0.5024$
Ours $\hat{M}$ $0.8534$ $0.6039$ $0.7481$
$M_{\text{AP}}$ $0.8358$ $0.5951$ $0.7429$
$M$ 0.8680 0.6494 0.7859
To evaluate the performance of the hair segmentation methods, we manually annotated hair regions in $150$ images from the test set. For baseline methods, we selected previous studies related to scalp image analysis  [44, 52, 23] and SAM [26]. Additionally, we performed an ablation study comparing our final segmentation result ($M$) with the intermediate results ($\hat{M}$ and $M_{\text{AP}}$).
Quantitative results. Table 1 reveals that our methods surpass the performance of existing hair segmentation techniques. In particular, the approach of combining the advantages of the two masks, $\hat{M}$ and $M_{\text{AP}}$ using the AND operator in $M$ showed the best performance. These results show the limitations of traditional computer vision techniques used in previous studies for image segmentation, revealing a lack of understanding in capturing the intricate patterns of hair and the scalp. Additionally, we observed that SAM, when used without specific guidance, was less effective for automatic image segmentation.
Qualitative results. As shown in Figure 5, our approach demonstrates effective hair segmentation with robustness to noise, providing clear and accurate hair segmentation compared to previous methods. Furthermore, it shows that $\hat{M}$ faces challenges in clearly capturing hair, and it exhibits robustness against noise such as dandruff. Conversely, $M_{\text{AP}}$ captures the hair well but is less robust to noise. Therefore, the combination of the two masks, $M$, demonstrates the mitigation of the drawbacks of each mask.
4.3 Scalp Disease and Severity Diagnosis
Figure 6: Image translation results with different generative models. The goal is to maintain the content of the source image while translating the scalp condition to the target image.
Figure 7: Image translation results using various mask guidance. Note that our approach is guided by $1-M$.
Table 2: Performance of scalp condition classification with various augmentation methods, denoted after “+” symbol, on the test set. The second column displays the overall macro-F1 score, while the third to last column shows the F1 scores for each severity level of three diseases.
Model F1 Dandruff Sebum Erythema
macro good mild moderate severe good mild moderate severe good mild moderate severe
DenseNet [21] $0.5819$ $0.7962$ $0.5140$ $0.5916$ $0.6138$ 0.5541 $0.6013$ $0.6412$ $0.0000$ 0.7758 $0.7287$ $0.5652$ $0.6010$
+ Gaussian Noise $0.5667$ $0.7804$ $0.4969$ $0.5662$ $0.5971$ $0.4712$ $0.5809$ $0.5946$ $0.0000$ $0.7509$ $0.7119$ $0.6142$ $0.6352$
+ AugMix [18] $0.5249$ $0.7886$ $0.5009$ $0.5890$ $0.0000$ $0.5042$ $0.5882$ $0.6341$ $0.0000$ $0.7430$ $0.7179$ $0.5956$ $0.5845$
+ DiffuseIT [27] $0.6079$ $0.8090$ $0.4824$ $0.6041$ $0.6500$ $0.5357$ $0.6126$ $0.6245$ $0.2024$ $0.7744$ $0.7395$ $0.6209$ $0.6388$
+ AGG [28] $0.6103$ $0.8106$ $0.4803$ $0.5908$ $0.6539$ $0.5175$ $0.5977$ $0.6119$ $0.2997$ $0.7713$ 0.7402 0.6291 $0.6206$
+ Ours 0.6358 0.8196 0.5407 0.6253 0.6652 $0.5362$ 0.6174 0.6414 0.4296 $0.7578$ $0.7339$ $0.6214$ 0.6412
EfficientFormerV2 [30] $0.5692$ $0.7954$ $0.4172$ $0.5983$ $0.6279$ $0.5257$ $0.5648$ $0.6283$ $0.0000$ $0.7722$ $0.7090$ $0.6227$ $0.5685$
+ Gaussian Noise $0.5618$ $0.7799$ $0.4774$ $0.5658$ $0.6326$ $0.4601$ $0.5850$ $0.5501$ $0.0000$ $0.7418$ $0.7141$ $0.5980$ $0.6372$
+ AugMix [18] $0.5770$ $0.7885$ $0.4940$ $0.5916$ $0.6347$ $0.5189$ $0.5927$ $0.6228$ $0.0000$ $0.7460$ $0.7238$ $0.6199$ $0.5904$
+ DiffuseIT [27] $0.5955$ $0.7981$ $0.4410$ $0.5977$ $0.6316$ $0.5263$ $0.5953$ $0.6057$ $0.2362$ $0.7658$ $0.7149$ $0.6121$ $0.6214$
+ AGG [28] $0.6098$ $0.8012$ $0.5087$ $0.6037$ $0.6262$ $0.5112$ $0.5830$ $0.6084$ $0.2999$ 0.7874 $0.7364$ $0.6239$ $0.6275$
+ Ours 0.6346 0.8066 0.5290 0.6190 0.6689 0.5345 0.6129 0.6324 0.4060 $0.7811$ 0.7382 0.6394 0.6476
Synthetic image generation. For the evaluation of DiffuseIT-M, we compared our model against DiffuseIT [27] and AGG [28] as baseline for the state-of-the-art image-to-image translation model. We have selected to employ the FID [19] and LPIPS [55] scores for fidelity evaluation. Comparative experiments were conducted using images from our augmentation dataset, with DiffuseIT and AGG serving as baseline models. Table 3 reveals that DiffuseIT-M outperforms other models in both metrics, indicating superior image fidelity. This high-quality image generation is attributed to our model’s effective implementation of mask guidance. As shown in Figure 6, we can observe that both DiffuseIT and AGG models fail to preserve the hair content information from the source image. Furthermore, these models tended to compromise overall information and were unable to transfer the semantic information. However, our model successfully preserved hair content information and transferred the semantic information.
Table 3: Quantitative analysis of image-to-image translation performance.
Model FID ($\downarrow$) LPIPS ($\downarrow$)
DiffuseIT [27] $138.42$ $0.4631$
AGG [28] $141.70$ $0.4915$
Ours 74.84 0.3528
Effect of mask guidance. We conducted experiments to examine the impact of mask guidance on hair information preservation during image translation. As illustrated in Figure 7, our method, guided by the mask $1-M$, effectively retains hair features while successfully transferring the semantic attributes of the target image onto the scalp. In contrast, using the reverse mask, $M$, leads to only minor alterations in scalp color from the target image, with a notable transfer of hair semantic information from the target. When no mask ($\mathbf{0}$) is applied, the translation results in minimal color change, failing to transfer conditions like dandruff from the target image. Conversely, with a full mask ($\mathbf{1}$), both hair and scalp features are subjected to changes. This differentiation in results highlights the importance of mask guidance in preserving specific image features, demonstrating the versatility of our approach in handling different translation objectives.
Scalp condition classification. To demonstrate the effectiveness of our augmentation method using generated images, we employed two different models as the classification backbone: DenseNet [21] as a CNN and EfficientFormerV2 [30] as a Transformer. We conducted a comparative analysis with other augmentation methods, including the addition of Gaussian noise, AugMix [18], DiffuseIT [27], and AGG [28]. The results are summarized in Table 2.
Our approach, specifically employing DiffuseIT-M, achieved the highest performance in both models. Notably, classifying the severe sebum class proved to be especially challenging when using non-generative augmentation methods. This difficulty arises primarily due to the extreme scarcity of samples for this class. The augmentation of the training dataset with generative models led to enhanced performance compared to the baseline, thanks to the effective style and content guidance from these models. Our model, in particular, exhibited superior accuracy compared to DiffuseIT and AGG, which struggled to preserve the essential information of the hair effectively. This underscores the significance of incorporating both the scalp style details and the hair content information in scalp disease classification.
4.4 Alopecia Severity Diagnosis
Table 4: Performance of alopecia severity classification.
Method Accuracy Weighted F1
Image-Based DenseNet [51] $0.6686$ $0.6283$
EfficientFormerV2 [12] $0.6728$ $0.6351$
Feature-Based Kim et al. [23] $0.5994$ $0.6062$
Ours 0.7460 0.6873
Figure 8: Distribution of features used for alopecia prediction. This figure illustrates the distribution of $\mathcal{T}$, representing hair thickness values, and $|B|$, indicating the count of bounding boxes in each image.
To diagnose the severity of alopecia, we utilized hair-related features such as hair thickness and count, as described in Section 3.2. These features were then employed to make predictions using a GradientBoosting model [15].
Our method was benchmarked against both image-based and traditional feature-based diagnostic approaches. For image-based severity prediction, we used DenseNet and EfficientFormerV2 models as baselines, which directly analyze the images to predict alopecia severity. Conversely, for feature-based prediction, we employed a method from [23], which involves generating masks through Otsu-based thresholding [34] and extracting features based on average hair thickness.
Since the data on alopecia severity is imbalanced, as shown in Figure 1, we selected to use both accuracy and weighted F1-score as evaluation metrics.
Alopecia severity classification. Our approach consistently outperforms other methods, including those based on image analysis, as shown in Table 4. This highlights the significance of hair thickness and counts as key diagnostic features for alopecia and validates the accuracy of our method in predicting these attributes precisely.
Feature distribution. We analyzed the distribution of input features across different alopecia severity levels. As illustrated in Figure 8, box plots for each feature show a consistent decrease from the good to severe categories. This trend indicates that our balanced sample set allowed each feature to contribute effectively to the severity prediction, enhancing the model’s accuracy.
Table 5: Comparative analysis of different prompting methods with SAM. The symbol “–” indicates scenarios where no prompt was used.
Positive Negative Pixel-F1 Jaccard Dice
$\hat{M}$ $1-\hat{M}$ $0.3786$ $0.1199$ $0.1891$
^$\hat{\mathcal{H}}_{\text{skel}}$ $1-\hat{M}$ $0.8210$ $0.5707$ $0.7215$
$\hat{C}$ $0.4966$ $0.2754$ $0.4148$
$\hat{C}$ $1-\hat{M}$ 0.8358 0.5951 0.7429
5 Ablation Studies
5.1 Automatic Prompting for SAM
We conducted an ablation study to evaluate our automatic prompting method with SAM. The study comprised four different approaches: 1) positive point prompts from $\hat{M}$ and negative point prompts from $1-\hat{M}$, 2) positive point prompts from ^$\hat{\mathcal{H}}_{\text{skel}}$ and negative point prompts from $1-\hat{M}$, 3) positive point prompts from $\hat{\mathcal{C}}$ without negative prompts, and 4) positive prompt points from $\hat{\mathcal{C}}$ and negative point prompts from $1-\hat{M}$ (Ours).
Quantitative results. Table 5 demonstrates that our method outperforms the others across all metrics. Notably, the absence of negative prompts significantly diminishes performance, underscoring their importance in the model.
Figure 9: Illustration of hair segmentation results using different prompt guides. “GT” represents segmentation guided by the ground truth mask, while $M_{\text{AP}}$ signifies our final prompting method with positive prompts from $\hat{\mathcal{C}}$ and negative prompts from $(1-\hat{M})$. “(+)(+)( + )” refers to positive prompts and “()(-)( - )” refers to negative prompts.
Qualitative results. Figure 9 showcases the efficacy of our point prompting method in hair segmentation. While using ^$\hat{\mathcal{H}}_{\text{skel}}$ for prompting improves over $\hat{M}$, it often results in unevenly distributed sampled points, leading to a noisier mask than that produced by our method.
5.2 Hair Preservation Efficacy in Generative Models
We conducted an analysis to understand how various image-to-image generative models impact hair segmentation. Specifically, we transformed 150 test images by applying a healthy target image (devoid of any scalp diseases) using three different generative models. We then applied our segmentation method to these images for a thorough quantitative evaluation.
Quantitative results. For a numerical evaluation of hair preservation, we generated images and then applied segmentation metrics like pixel F1, Jaccard index, and Dice score. Table 6 showcases the results, highlighting that DiffuseIT-M (denoted as Ours ($M$)) excels in preserving hair details, thereby affirming its efficacy in retaining essential content from the source images. Moreover, we examined the segmentation outcomes when using GT masks (denoted as Ours (GT)) alongside our model. This comparison illustrated that improving mask quality significantly enhances the preservation of hair features in the generated images, providing insights into how mask quality impacts hair content retention in generative models.
Qualitative results. We compared hair preservation in images generated by our model with those from other models. Our model used GT masks as guides during the generation process. Figure 10 shows the hair segmentation results for images generated by our model alongside the GT masks. The comparison demonstrates that our model effectively retains hair content from source images, unlike DiffuseIT and AGG, which struggled to maintain comprehensive hair details. In particular, AGG tended to preserve only dominant hair features, omitting finer details.
Table 6: Comparative analysis of hair segmentation across different generative methods.
Model Pixel-F1 Jaccard Dice
DiffuseIT $0.7106$ $0.3619$ $0.5153$
AGG $0.7286$ $0.3785$ $0.5403$
Ours ($M$) $0.7625$ $0.4485$ $0.6067$
Ours (GT) 0.7812 0.4840 0.6416
Figure 10: Comparative qualitative results of hair segmentation across different generative models.
Table 7: Ablation study of impact analysis of features in the alopecia severity prediction. $HC$ is the number of Hair estimation from Section 3, $\mathcal{T}$ is the list of hair thickness estimations and $|B|$ is the number of bounding boxes in an image. The “ $\tilde{\mathcal{T}}$ ” symbol denotes the median of $\mathcal{T}$, while the “ $\overline{\mathcal{T}}$ ” symbol signifies the mean of $\mathcal{T}$. An $\ast$” symbol signifies that outliers were removed. The order in which features are added reflects feature importance in the final model.
$HC$ $\tilde{\mathcal{T}}^{\ast}$ $\tilde{\mathcal{T}}$ $|B|$ $\overline{\mathcal{T}}$ $\overline{\mathcal{T}}^{\ast}$ Accuracy F1 (Weighted)
$0.7233$ $0.6109$
$0.7369$ $0.6650$
$0.7390$ $0.6712$
$0.7407$ $0.6761$
$0.7436$ $0.6783$
0.7460 0.6873
5.3 Feature Selection for Severity Estimation
We identified key features for alopecia severity prediction, focusing on hair thickness and count. For thickness, we analyzed the length of the list $B$ and computed the mean and median values of $\mathcal{T}$, excluding outliers identified using the interquartile range (IQR). We also considered the mean and median of the entire $\mathcal{T}$. Additionally, the number of line segments detected by the hair count estimator was used as a supplementary feature.
Feature importance. An analysis using the GradientBoosting model (Table 7) revealed that the best prediction accuracy was achieved when all features were combined. This finding underscores the significance of each extracted feature in our model.
6 Conclusion
In this work, we introduced ScalpVision, diagnostic system designed for a complete evaluation of scalp health. This system harnesses label-free learning combined with a foundational model to create precise hair segmentation masks. Additionally, ScalpVision employs diffusion-based image augmentation for accurate scalp condition diagnosis and leverages estimations of hair thickness and count to assess alopecia severity. We envision ScalpVision as a significant step towards a generalized diagnostic system for dermatological applications, particularly in simplifying the challenges associated with complex labeling in microscopic scalp imaging.
7 Limitation
There are several limitations of ScalpVision that would make for promising avenues of future work, which mainly steams from dataset constraints. The limited labels in the dataset pose challenges in accurately diagnosing scalp and alopecia conditions. For example, while follicle count is crucial for alopecia diagnosis, we used hair count as a proxy due to data limitations. However, not all visible hairs in an image necessarily originate from follicles within the same image, making this a less than ideal substitute. Additionally, we assumed that scalp images capture a consistent area. In reality, slight variations in the distance from which images are captured can lead to differences in perceived hair thickness. Addressing these discrepancies requires further research into standardizing image capture methods or developing models robust to such variations.
Acknowledge
This work was partly supported by an IITP grant (No. RS-2020-II201361, RS-2022-00155966) and NRF grant (No. RS-2024-00354218) funded by the Korea government (MSIT).
Supplementary Material
This supplementary material provides additional details and extended results that complement the main paper. Included in this supplementary material are the following sections:
  • Section A delves into detailed explanations of various medical scalp diseases.
  • Section B offers a comprehensive look at our experimental setup and provides supplementary information.
  • Section C introduces pseudo training set for training the hair segmentation model.
  • Section D presents further insights into our approach to scalp disease classification.
  • Section E extends the discussion on the selection and analysis of features used in predicting the severity of alopecia.
  • Section F shows additional qualitative results of DiffuseIT-M.
Appendix A Detailed Overview of Scalp Diseases
The dataset from AI-Hub [^3] https://aihub.or.kr categorizes scalp images into three primary conditions: dandruff, excess sebum, and erythema.
Dandruff, also referred to as a milder manifestation of seborrheic dermatitis, is characterized by the non-inflammatory exfoliation of dead epidermal cells from the scalp. While it can induce mild itching, it generally does not precipitate erythema or the formation of scabs [4].
Hyperseborrhea, the excessive production of sebum, represents a common aesthetic concern, manifested through the secretion of excess oil from hypertrophic sebaceous glands. This condition results in a shiny, oily skin appearance. Although sebum plays a crucial role in maintaining skin hydration and its protective barrier, its excessive secretion can lead to various dermatological issues. One such issue is the formation of sebum plugs, which are small, yellowish, or pale bumps that appear on the skin [42].
Scalp erythema, also known as red scalp, is characterized by widespread redness across the scalp. It can arise from several conditions, including psoriasis, seborrheic dermatitis, contact dermatitis, diffuse lichen planopilaris, dermatomyositis, and scalp rosacea [49].
Appendix B Implementation Details
Hair segmentation. For our heuristic-driven hair segmentation, we utilized the U2-Net [39] official source code to train $3{,}000$ pseudo image-label pairs. The training parameters included a batch size of $32$, a constant learning rate of $0.001$, $100$ training epochs, and the Adam optimizer. The output from U2-Net was binarized to obtain $\hat{M}$, using a threshold of $0.5$.
Since other previous studies, excluding SAM [26], do not have an official codebase, we implemented their approaches based on descriptions in their papers, using OpenCV. For instance, [44] used contrast stretching and binary thresholding to derive the hair mask. [52] applied morphological operations, while [23] used Otsu’s method [34] for hair mask acquisition. In evaluating SAM, the mask with the highest Intersection over Union (IoU) score was selected as the final prediction.
Image augmentation. For DiffuseIT-M, detailed in Section 3.3 of the main paper, we construct a loss function for image translation while preserving hair content. This function is defined as follows:
$\begin{gathered}\ell_{total}(x;x_{src},x_{trg},M)=\lambda_{1}\ell_{style}+\lambda_{2}\ell_{content}\\ +\lambda_{3}\ell_{mask}+\lambda_{4}\ell_{sem}+\lambda_{5}\ell_{rng}.\end{gathered}$ (18)
To incorporate the semantic information of the target image, we establish the style loss function, $\ell_{sty}$. This function leverages the [CLS] token from the last layer of DINO-ViT [6]. Denoting the [CLS] tokens as $\mathbf{c}$, the style loss is expressed as:
$\begin{gathered}\ell_{sty}(x_{trg},\hat{x}_{0}(x_{t}))=||\mathbf{c}(x_{trg})-\mathbf{c}(\bar{x})||_{2}\\ +\lambda_{mse}||x_{trg}-\hat{x}_{0}(x_{t})||_{2},\end{gathered}$ (19)
where $\lambda_{mse}$ is set to $3{,}000$, and the weight for $\ell_{style}$, $\lambda_{1}$, is set to $2{,}000$.
The content loss, $\ell_{content}$, is designed to preserve the structure of source images. Let $k^{l}_{i}(x)$ represent the $i$-th key extracted from the $l$-th multi-head self-attention layer in DINO-ViT for image $x$. The content loss is then defined as:
$\ell_{content}=\lambda_{sim}\ell_{sim}(x_{src},\hat{x}_{0}(x_{t}))+\lambda_{con}\ell_{con}(x_{src},\hat{x}_{0}(x_{t})),$ (20)
where the similarity loss, $\ell_{sim}$, and the content loss, $\ell_{con}$, are
$\ell_{sim}(x_{src},\hat{x}_{0}(x_{t}))=||\text{cos}_{ij}(x_{src}),\text{cos}_{ij}(\hat{x}_{0}(x_{t}))||_{2},$ (21)
$\ell_{con}(x_{src},\hat{x}_{0}(x_{t}))=\text{infoNCE}(k^{l}_{i}(x_{src}),k^{l}_{i}(\hat{x}_{0}(x_{t}))),$ (22)
with $cos_{ij}(x)$ representing the cosine distance between $k^{l}_{i}(x)$ and $k^{l}_{j}(x)$. The weights $\lambda_{sim}$ and $\lambda_{con}$ are set to $1{,}000$ and $200$, respectively. Additionally, weights $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to $1,000$, $100$, and $200$, respectively. During our experiment, the model is configured to generate images with a resolution of $256\times 256$ pixels, utilizing a denoising step of $1{,}000$.
For implementation previous studies, we have performed the official source code of both DiffuseIT [27] and AGG [28], conducting experiments with the provided hyperparameters.
For the implementation of previous studies, we used the official source code of DiffuseIT [27] and AGG [28], conducting experiments with their provided hyperparameters. Non-generative image augmentations were performed using AugMix [18] through the torchvision library and Gaussian noise augmentation via the PyTorch library.
Figure 11: Examples of pseudo images and their corresponding masks for hair segmentation.
Scalp disease and severity diagnosis. For the classification task, we fine-tuned two models: DenseNet169 [21] (CNN-based) and EfficientFormerV2 [30] (Transformer-based), using pre-trained weights from the timm library [50]. Fine-tuning involved a batch size of $128$, a learning rate of $0.0001$ with a CosineAnnealingWarmRestarts scheduler [31], $50$ training epochs, and the AdamW optimizer [32].
Alopecia severity prediction. To predict the severity of alopecia, we focused on extracting hair-related features, specifically hair thickness and count. For this, we employed the OpenCV library, utilizing Non-Maximum Suppression (NMS), Morphology operation, and Hough transformation techniques. These features were then used to train a prediction model with the GradientBoosting Classifier, a component of the Scikit-learn library [5]. The model’s training parameters were set as follows: a cross-entropy loss function, a learning rate of $0.01$, a maximum depth of $5$, $3{,}000$ estimators, and a subsample ratio of $1$. All other parameters remained at their default settings in Scikit-learn.
In comparison, for the image-based models, we trained DenseNet169 and EfficientFormerV2 using similar hyperparameters as those used for scalp disease classification. The key differences were the employment of the cross-entropy loss function and the use of the Adam optimizer for these models.
Appendix C Pseudo Image and Mask Visualization
To create a diverse pseudo training set, we extracted scalp patch images from areas without hair in nine different scalp images. Each image represented a unique disease at a specific severity level. As illustrated in Figure 11, we introduced a variety of hair types by inserting straight and curved lines in blue, brown, black, and white colors, each with differing thicknesses. To simulate common scalp noise, such as dandruff, white circular elements were added to the pseudo images. Our codebase contains further details on this process.
Appendix D Scalp Disease and Severity Classification
This section outlines our data augmentation approach for classifying scalp diseases and their severities and presents additional experimental results.
Algorithm 2 Calculation of sampling ratios for each disease
Input: $severities$: A collection of records for each severity. Assume that each element of $severities$ is classified as 0 (good), 1 (mild), 2 (moderate), and 3 (severe).
Output: $ratios$: Sampling ratio among four severity levels.
1:  $\epsilon\leftarrow 1\times 10^{-9}$
[…]
3:  for each $sevLevel$ in $[0,1,2,3]$ do
4:     $sevCounts[sevLevel]$leftarrow 00
5:  end for
6:  for each $severity$ in $severities$ do
7:     $sevCounts[severity]$leftarrow $sevCounts[severity]+1$
8:  end for
[…]
10:  for each $sevLevel$ in $[0,1,2,3]$ do
11:     $invCounts[sevLevel]$leftarrow $1/(sevCounts[sevLevel]+\epsilon)$
12:  end for
13:  $normFactor$leftarrow $\sum_{sevLevel=0}^{3}invCounts[sevLevel]$
14:  $ratios$leftarrow [][][ ]
15:  for each $sevLevel$ in $[0,1,2,3]$ do
16:     $ratio\leftarrow invCounts[sevLevel]/normFactor$
17:     $ratios[sevLevel]$leftarrow $ratio$
18:  end for
19:  return $ratios$
D.1 Data Augmentation Strategy
Figure 12: Data distribution of AI-Hub scalp dataset before and after the application of augmentation techniques.
Addressing the issue of data imbalance in the dataset, we implemented a strategy to translate randomly selected images into classes with fewer samples. We used random selection for the source images and weighted sampling for target images, where the likelihood of choosing an image was inversely proportional to the number of samples in its severity class. This method favored the selection of underrepresented classes. The algorithm to calculate these sampling weights is detailed in Algorithm 2.
Figure 12 displays the data distribution before and after this augmentation. The post-augmentation distribution shows a more balanced representation across various classes, especially a rise in the severe category and a reduction in the good category. This balanced dataset played a crucial role in enhancing our model’s performance by providing an even distribution of the training samples. The same augmentation strategy was applied across all diffusion-based methods.
D.2 Impact of Backbone Model
We assessed the performance of scalp disease classification using different pretrained backbone models. Our evaluation metric was the F1 macro score, which we also used to compare our results with existing augmentation methods, including those utilizing DiffuseIT and AGG. The experiments involved two CNN-based models (ResNet [16] and ResNeXt [51]) and two Transformer-based models (ViT [12] and EfficientFormer [29]), all maintaining consistent hyperparameters as described in Section 11.
As presented in Table 8, our approach achieved superior classification performance across all backbone models. Moreover, models augmented using generative methods consistently outperformed those using baseline augmentation, validating the effectiveness of our proposed data augmentation strategy.
Table 8: Quantitative comparison of classification performance across different backbone models. “Baseline” refers to performance without any augmentation methods, while results following the “+” symbol indicate the use of various augmentation methods. Values in the table represent the macro-F1 scores.
Method CNN Transformer
ResNet ResNeXt EfficientFormer ViT
Baseline $0.5709$ $0.6186$ $0.5300$ $0.4934$
+ Gaussian Noise $0.4304$ $0.5663$ $0.5742$ $0.4306$
+ AugMix $0.4186$ $0.5247$ $0.5348$ $0.4747$
+ DiffuseIT $0.6064$ $0.6107$ $0.6120$ $0.5767$
+ AGG $0.5978$ $0.6251$ $0.6104$ $0.5771$
+ Ours 0.6128 0.6292 0.6170 0.5861
Figure 13: Qualitative results of hair count estimation. The displayed lines on the masks represent the line segments identified through the Hough Transformation.
Figure 14: Qualitative results and estimated features of hair thickness estimation. The red dots on the image indicate the intersection points of hair strands and perpendicular line of hair direction, while the green lines represents the perpendicular lines. Each caption represents severity of alopecia.
Appendix E Alopecia Severity Prediction
E.1 Feature Selection for Severity Estimation
We identified key features for alopecia severity prediction, focusing on hair thickness and count. For thickness, we analyzed the length of list $B$ and computed the mean and median values of $\mathcal{T}$, excluding outliers identified using the interquartile range (IQR). We also considered the mean and median of the entire $\mathcal{T}$. Additionally, the number of line segments detected by the hair count estimator was used as a supplementary feature.
E.2 Visualization of Severity Features
Our method estimates hair thickness using mask images. Figure 13 showcases the qualitative results of our multi-scale Hough transform, which effectively distinguishes individual hairlines in the mask images. In Figure 14, we demonstrate thickness estimation across various scenarios for four alopecia severity classes: good, mild, moderate, and severe. Our segmentation successfully identified different hair thicknesses and distinguished hair from severe scalp conditions like dandruff or erythema. Notably, as alopecia severity increased, the feature values related to hair thickness decreased, validating the effectiveness of our approach.
Appendix F Qualitative Results of DIffuseIT-M
Figure 15 presents qualitative results of our model’s ability to translate images across multiple labels while maintaining hair information. The results indicate that our model successfully translates various disease features from target images, effectively preserving hair representation.
Figure 15: Qualitative results of DiffuseIT-M. This figure illustrates the results for various scalp disease conditions, with severity levels indicated as 0 (good), 1 (mild), 2 (moderate), and 3 (severe). Scalp diseases are color-coded for clarity: blue represents dandruff, green signifies excess sebum, and red denotes erythema.