### v3

# Scalp Diagnostic System With Label-Free Segmentation and Training-Free Image Translation

Abstract
[…]
Keywords:
Scalp Disease Diagnosis Generative Data Augmentation.
1 Introduction
Scalp disorders are a widespread concern, with nearly $90\%$ of adults in the U.S. experiencing some form of condition [8]. Left unchecked, even seemingly mild scalp ailments can escalate into more serious outcomes, such as alopecia, underscoring the importance of timely intervention. Consequently, early diagnosis is crucial for preventing the progression of scalp-related diseases [23, 24], highlighting the need for advanced diagnostic approaches that are both efficient and accessible. Recognizing the importance of early detection, numerous studies have explored scalp disease diagnosis using microscopic scalp imagery [5, 14, 27].
Nevertheless, effectively diagnosing scalp disorders relies heavily on measuring critical features such as hair count and thickness, which demand precise hair segmentation. However, generating pixel-level hair annotations is costly and time-consuming, and no publicly available dataset provides such segmentation labels. The only major resource, AI-Hub [1], offers classification labels for scalp conditions but lacks segmentation annotations (see Section 3.1). Moreover, like many scalp image datasets, it suffers from data imbalance, especially for severe conditions, making it challenging to develop robust models.
To overcome these limitations, we propose ScalpVision, a comprehensive system for the in-depth assessment of scalp health. First, we achieve label-free hair segmentation by combining a naive segmentation model – trained on synthetic image-label pairs – with an automatic prompting module for the Segment Anything Model (SAM) [15], systematically generating positive and negative point prompts to enable accurate hair masks without manual labeling. Building on these masks, we then introduce DiffuseIT-M, a diffusion-based image-to-image translation framework that preserves hair details while altering scalp conditions. By generating diverse training samples, our method effectively mitigates data imbalance, ultimately leading to enhanced diagnostic performance for scalp diseases.
Figure 1: ScalpVision pipeline overview: II is the original image, model $S$ generates the hair segmentation mask $\hat{M}$ using a pseudo-training set, $M_{\text{AP}}$ is the SAM-produced mask, and $M$ is the combined hair segmentation mask. The “Automatic Prompt” for refining segmentation comes from $\hat{M}$. $x_{src}$ and $x_{trg}$ are the source and target images, with $M$ as the mask image of $x_{src}$. The weighted image sum is denoted by $\odot$ and $D$ stands for DINO-ViT [4].
2 Method
As illustrated in Figure 1, central to ScalpVision is a hair segmentation module (Section 2.1) and an image translation module for generating diverse scalp images to augment training datasets for scalp condition classification (Section 2.2).
2.1 Label-Free Hair Segmentation
For the precise diagnosis of scalp conditions, our initial step involves segmenting hair within microscopic scalp images. However, since most scalp condition datasets lack segmentation labels, supervised learning methods are not feasible.
Heuristic-driven pseudo-labeling. To address the absence of hair segmentation, we first generate pseudo labels for training our segmentation model (S as shown in Figure 1) using prior knowledge. With the intuition that the hair on the microscopic scalp images follows either a linear function or a power function, we generate synthetic images to effectively guide the model to learn hair patterns on the scalp images. For each disease condition, we randomly select one image representing each distinct severity level, extract three smaller patches from regions of the scalp with no visible hair, and draw curves to simulate hair patterns. Additionally, to simulate dandruff noise, circular white shapes are added to these patches but are not indicated in the pseudo masks, thus training the model to interpret them as noise. Examples of the pseudo images and masks are provided in the Appendix 0.C. We generate 3,0003{,}000 pseudo-images and corresponding pseudo mask labels, using them to train the U2-Net [25] which generates the binary mask, $\hat{M}=\hat{M}(i,j)\in\{0,1\}^{H\times W}$, where $H$ and $W$ are the height and width of the image, and $i\in[1,H]$, $j\in[1,W]$ denote pixel coordinates.
Input: Mask $\hat{M}$, bounding box size $n$, cross-shaped structuring element kernel
Output: Representative hair points from mask $\hat{\mathcal{C}}$
1
[…]
3
4while $\mathcal{H}_{\text{copy}}\neq 0$ do
5  Eroded $\leftarrow\texttt{MORPHOLOGY}\_\texttt{ERODE}(\mathcal{H}_{\text{copy}},\textit{kernel})$
6  $\textit{Dilated}\leftarrow\texttt{MORPHOLOGY}\_\texttt{DILATE}(\textit{Eroded},\textit{kernel})$
7  $\hat{K}\leftarrow\mathcal{H}_{\text{copy}}-\textit{Dilated}\,$; $\hat{\mathcal{H}}_{\text{skel}}\leftarrow\hat{\mathcal{H}}_{\text{skel}}\lor\hat{K}\,$; $\mathcal{H}_{\text{copy}}\leftarrow\textit{Eroded}$
8
9
10foreach $(x,y)\in\mathcal{H}_{\text{skel}}$ do
[…]
12
13
[…]
15foreach $(x_{1},y_{1},x_{2},y_{2})\in\hat{B}$ do
[…]
17
18
19return $\hat{\mathcal{C}}$
Algorithm 1 Extraction of representative points from mask
Automatic prompting for SAM. To refine the hair segmentation mask $\hat{M}$, we utilize the foundation segmentation model, SAM [15], employing a point-prompting method to differentiate hair from scalp without additional training. However, selecting random points from $\hat{M}$ for positive prompts often led to suboptimal masks, mainly due to points near the edges of $\hat{M}$ confusing the SAM. Furthermore, the intrinsic randomness occasionally caused sampled points to coalesce within a confined region, thereby leading the SAM to segment only a limited subset of hairs. To address these issues, we developed an automatic prompting method, shown in Algorithm 1, that uniformly samples across $\hat{M}$ and uses the coarse segmentation mask $\hat{M}$ to guide the SAM with high confidence.
To extract the distinct features of the hair, we compute the skeletonized mask, $\hat{\mathcal{H}}_{\text{skel}}\in\{0,1\}^{H\times W}$, using morphological erosion and dilation following [34]. Then, we generate bounding boxes around each pixel in $\hat{\mathcal{H}}_{\text{skel}}$ with size $n\times n$ where we set $n=10$. These boxes undergo non-maximum suppression (NMS) to filter out the bounding boxes, denoted as $\hat{B}=\{\hat{b}_{j}\}_{j=1}^{k}$, where each box is defined by coordinates $(x_{\text{min}},y_{\text{min}},x_{\text{max}},y_{\text{max}})$. Following this, the mean points of the hair pixels, $\hat{\mathcal{C}}=\{\hat{c}_{j}\}_{j=1}^{k}$, in each bounding box $\hat{B}$ can be determined. For each $\hat{b}_{j}=(x_{1},y_{1},x_{2},y_{2})$, the mean point $\hat{c}_{j}=(\bar{x},\bar{y})$ is given by:
$\bar{x}=\frac{\sum_{i}\sum_{j}i\cdot\hat{\mathcal{H}}(i,j)}{\sum_{i}\sum_{j}\hat{\mathcal{H}}(i,j)},\quad\bar{y}=\frac{\sum_{i}\sum_{j}j\cdot\hat{\mathcal{H}}(i,j)}{\sum_{i}\sum_{j}\hat{\mathcal{H}}(i,j)}$ (1)
where the summation is over all $i\in[x_{1},x_{2}]$ and $j\in[y_{1},y_{2}]$.
Subsequently, we select positive point prompts for SAM from the calculated mean points $\hat{\mathcal{C}}$. For the negative point prompts, we utilize the inverse of the initial mask, specifically $1-\hat{M}$. These prompts, automatically generated, guide SAM in generating the binary segmentation mask, $M_{\text{AP}}\in\{0,1\}^{H\times W}$.
Mask ensemble. $M_{\text{AP}}$ and $\hat{M}$ complement each other with strengths and weaknesses. $\hat{M}$ is robust against noise like dandruff as it was trained using simulated noise. Meanwhile, $M_{\text{AP}}$, benefiting from SAM’s superior edge detection, excels in constructing a clear boundary between hair and scalp. Therefore, to make a robust hair mask, the final binary mask, $M$, is derived from $\hat{M}$ and $M_{\text{AP}}$ with the logical AND operation ($M=\hat{M}\land M_{\text{AP}}$), followed by a noisy region removal post-processing step with connected-component analysis.
2.2 Scalp Condition Classification
Accurately classifying scalp disease severity from microscopic images is difficult due to the rarity of extreme cases. To address this, we introduce DiffuseIT-M, a diffusion-based image translation model with mask guidance that transforms a source image into various scalp conditions while preserving hair content. Building on DiffuseIT [16] and incorporating an image editing technique inspired by blended diffusion [2], DiffuseIT-M enables robust augmentation of underrepresented classes for improved classification.
Image translation with mask guidance. To facilitate the transfer of scalp disease characteristics while preserving hair features in our model, we utilize a comprehensive loss function, $\ell_{total}$, that guides the reverse process and is composed of five distinct loss components. These components consider the source image ($x_{src}$), the target image ($x_{trg}$), and the hair mask ($M$) as inputs. The combined loss function is defined as:
$\begin{gathered}\ell_{total}\left(\scriptstyle{x;x_{src},x_{trg},M}\right)=\lambda_{1}\ell_{style}+\lambda_{2}\ell_{content}+\lambda_{3}\ell_{mask}+\lambda_{4}\ell_{sem}+\lambda_{5}\ell_{rng},\end{gathered}$ (2)
where $\lambda_{i\in[1,5]}$ denotes the weights assigned to each of these loss functions.
For $\ell_{style}$ and $\ell_{content}$, we utilize the style and content loss functions from DiffuseIT, with further details provided in the Appendix 0.B. We employ the [CLS] token matching loss using DINO-ViT [4] to reflect semantic information in $x_{trg}$ and use keys of multi-head self-attention layers to preserve the content of $x_{src}$. Additionally, to ensure hair preservation while translating scalp styles, we construct a mask preservation loss function as:
[…]
where LPIPS denotes the learned perceptual image patch similarity metric [33] and $\hat{x}_{0}(x_{t})$ is the estimation of the cleaned image derived from the sample $x_{t}$:
$\hat{x}_{0}(x_{t})=\frac{x_{t}}{\sqrt{\bar{\alpha}_{t}}}-\frac{\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$ (4)
We also include two additional losses: $\ell_{rng}$, representing the squared spherical distance as proposed in [6], and $\ell_{sem}$, indicating the semantic divergence loss as outlined in [16]. Using this composite loss function, $\ell_{total}$, we guide the generation of the next sample step, $x_{t-1}$. To preserve hair details, we apply a masking approach:
$x_{t-1}\leftarrow x_{t}\odot M+\bigl{[}\hat{x}_{0}(x_{t})-\nabla{x_{t}}\ell_{total}\bigl{(}\hat{x}_{0}(x_{t})\bigr{)}\bigr{]}\odot(1-M).$ (5)
This method allows scalp style translation without extra training.
Classification strategy. Using DiffuseIT-M, we augment our training set by translating randomly chosen images into higher severity levels via weighted sampling, where selection probability is inversely proportional to the class’s size. The augmentation strategy and the corresponding results are provided in the Appendix 0.D. We fine-tune a pretrained backbone with four MLP heads, each tied to a specific loss. One head detects the presence of scalp diseases (dandruff, excess sebum, erythema), while the other three classify their severities (good, mild, moderate, severe). Our objective, $\ell_{cls}$, is the sum of four losses: $\ell_{dis}$ (binary cross-entropy for disease presence), and $\ell_{dand}$, $\ell_{seb}$, $\ell_{ery}$(cross-entropy for severity classification). This design enables simultaneous disease detection and severity assessment.
Figure 2: Data distribution of different severity within each scalp condition.
3 Experiments
3.1 Dataset
Despite the lack of publicly available datasets, we accessed a specialized dataset from AI-Hub [1] for classifying the severity of scalp dermatologic conditions. $\dagger$$\dagger\daggerThis$ dataset is provided by ‘The Open AI Dataset Project (AI-Hub, S. Korea)’ and is exempt from IRB approval as it does not contain any information that can identify individuals. The dataset is publicly accessible at https://aihub.or.kr. The dataset comprises 95,91095{,}910 images with a resolution of $640\times 480$ pixels from 20,00020{,}000 patients. It is split into 72,34272{,}342 training and 23,56823{,}568 test images, with 21,70321{,}703 from the training set used for validation. Dermatologists labeled each image for dandruff, excess sebum, and erythema, categorizing severity as good, mild, moderate, or severe. The dataset is heavily skewed toward good and mild cases (Fig. 2) and lacks segmentation labels. Thus, we manually annotated hair regions in $150$ test images to evaluate our hair segmentation methods.
Figure 3: Comparison of various segmentation methods on hair. “GT” represents the mask images for which we have manually annotated the pixel segmentation. Note that $\hat{M}$, $M_{\text{AP}}$ and $M$ are proposed in Section 2.1.
Table 1: Performance of hair segmentation on the test set.
Approach Pixel-F1 Jaccard Dice
Shih et al. [28] $0.706$ $0.348$ $0.512$
Yue et al. [32] $0.794$ $0.493$ $0.654$
Kim et al. [13] $0.815$ $0.561$ $0.708$
SAM [15] $0.503$ $0.361$ $0.502$
Ours $\hat{M}$ $0.853$ $0.604$ $0.748$
$M_{\text{AP}}$ $0.836$ $0.595$ $0.743$
$M$ 0.868 0.649 0.786
Table 2: Quantitative analysis of image-to-image translation.
Model FID ($\downarrow$) LPIPS ($\downarrow$)
DiffuseIT $138.42$ $0.463$
AGG $141.70$ $0.492$
Ours 74.84 0.353
3.2 Hair Segmentation
Because narrow hairs are difficult to segment, many existing unsupervised methods still rely on traditional computer vision for hair-specific challenges. Accordingly, we compare prior scalp segmentation approaches [28, 32, 13] as baselines and also evaluate the foundational model SAM [15]. In addition, we conduct an ablation study on our final mask $M$ and its intermediate versions $\hat{M}$ and $M_{\mathrm{AP}}$.
Quantitative results. Table 2 reveals that our methods surpass the performance of existing hair segmentation techniques. In particular, the approach of combining the advantages of the two masks, $\hat{M}$ and $M_{\text{AP}}$ using the logical AND operator in $M$ showed the best performance. These results show the limitations of traditional computer vision techniques used in previous studies for image segmentation, revealing a lack of understanding in capturing the intricate patterns of hair and the scalp. Additionally, SAM was less effective for automatic segmentation when used without specific guidance.
Qualitative results. As shown in Figure 3, our approach demonstrates effective hair segmentation with robustness to noise, providing clear and accurate hair segmentation compared to previous methods. Furthermore, it shows that $\hat{M}$ faces challenges in clearly capturing hair, and it exhibits robustness against noise such as dandruff. Conversely, $M_{\text{AP}}$ captures the hair well but is less robust to noise. Therefore, the combination of the two masks, $M$, demonstrates the mitigation of the drawbacks of each mask.
3.3 Synthetic Image Generation
For the evaluation of DiffuseIT-M, we compared our model against DiffuseIT [16] and AGG [17] as baselines for the image-to-image translation model. This evaluation demonstrates that our model not only achieves high fidelity in image-to-image translation but also effectively preserves the desired hair details.
Figure 4: Image translation results with different generative models, where the goal is to preserve source hairlines while changing the scalp.
Figure 5: Image translation results using various mask guidance. Note that our approach is guided by 1M1-M.
Quantitative results. We have selected to employ the FID [11] and LPIPS [33] scores for fidelity evaluation using images from our augmentation dataset, with DiffuseIT and AGG serving as baseline models. Table 2 reveals that DiffuseIT-M outperforms other models in both metrics, indicating superior image fidelity. This high-quality image generation is attributed to our model’s effective implementation of mask guidance.
Qualitative results. Figure 5 shows that both DiffuseIT and AGG models fail to preserve the hair content information from the source image. Furthermore, these models tended to compromise overall information and were unable to transfer the semantic information. However, our model successfully preserved hair content information and transferred the semantic information.
Effect of mask guidance. We conducted experiments to examine the impact of mask guidance on hair information preservation during image translation. As illustrated in Figure 5, our method, guided by the mask 1M1-M, effectively retains hair features while successfully transferring the semantic attributes of the target image onto the scalp. In contrast, using the reverse mask, $M$, leads to only minor alterations in scalp color from the target image, with a notable transfer of hair semantic information from the target. When no mask ($\mathbf{0}$) is applied, the translation results in minimal color change, failing to transfer conditions like dandruff from the target. Conversely, with a full mask ($\mathbf{1}$), both hair and scalp features are subjected to changes. This differentiation in results highlights the importance of mask guidance in preserving specific image features, demonstrating the versatility of our approach in handling different translation objectives.
Table 3: Performance of scalp condition classification with various augmentation methods, denoted after “+” symbol, on the test set. The second column displays the overall macro-F1 score, while the columns from the third onward show the F1 scores for each severity level of the three diseases.
Model F1 Dandruff Sebum Erythema
macro good mild moderate severe good mild moderate severe good mild moderate severe
DenseNet $0.582$ $0.796$ $0.514$ $0.592$ $0.614$ 0.554 $0.601$ $0.641$ $0.000$ 0.776 $0.729$ $0.565$ $0.601$
+ Gaussian Noise $0.567$ $0.780$ $0.497$ $0.566$ $0.597$ $0.471$ $0.581$ $0.595$ $0.000$ $0.751$ $0.712$ $0.614$ $0.635$
+ AugMix $0.525$ $0.789$ $0.501$ $0.589$ $0.000$ $0.504$ $0.588$ $0.634$ $0.000$ $0.743$ $0.718$ $0.596$ $0.585$
+ DiffuseIT $0.608$ $0.809$ $0.482$ $0.604$ $0.650$ $0.536$ $0.613$ $0.625$ $0.202$ $0.774$ $0.740$ $0.621$ $0.639$
+ AGG $0.610$ $0.811$ $0.480$ $0.591$ $0.654$ $0.518$ $0.598$ $0.612$ $0.300$ $0.771$ 0.740 0.629 $0.621$
+ Ours 0.636 0.820 0.541 0.625 0.665 $0.536$ 0.617 0.641 0.430 $0.758$ $0.734$ $0.621$ 0.641
EfficientFormerV2 $0.569$ $0.795$ $0.417$ $0.598$ $0.628$ $0.526$ $0.565$ $0.628$ $0.000$ $0.772$ $0.709$ $0.623$ $0.569$
+ Gaussian Noise $0.562$ $0.780$ $0.477$ $0.566$ $0.633$ $0.460$ $0.585$ $0.550$ $0.000$ $0.742$ $0.714$ $0.598$ $0.637$
+ AugMix $0.577$ $0.789$ $0.494$ $0.592$ $0.635$ $0.519$ $0.593$ $0.623$ $0.000$ $0.746$ $0.724$ $0.620$ $0.590$
+ DiffuseIT $0.596$ $0.798$ $0.441$ $0.598$ $0.632$ $0.526$ $0.595$ $0.606$ $0.236$ $0.766$ $0.715$ $0.612$ $0.621$
+ AGG $0.610$ $0.801$ $0.509$ $0.604$ $0.626$ $0.511$ $0.583$ $0.608$ $0.300$ 0.787 $0.736$ $0.624$ $0.628$
+ Ours 0.635 0.807 0.529 0.619 0.669 0.535 0.613 0.632 0.406 $0.781$ 0.738 0.639 0.648
3.4 Scalp Condition Classification
To demonstrate the effectiveness of our augmentation method using generated images, we employed two different models as the classification backbone: DenseNet  [12] as a CNN and EfficientFormerV2 [18] as a Transformer. As summarized in Table 3, we compared only those augmentation methods that preserve a one-to-one correspondence between each image and its original, unambiguous set of condition labels – Gaussian noise, AugMix [10], DiffuseIT [16], and AGG [17]. Our approach, which specifically employs DiffuseIT-M, achieved the highest performance in both models. Notably, classifying the severe sebum class proved to be especially challenging when using non-generative augmentation methods. This difficulty arises primarily due to the extreme scarcity of samples for this class. The augmentation of the training dataset with generative models led to enhanced performance compared to the baseline. Our model, in particular, exhibited superior accuracy compared to DiffuseIT and AGG, which struggled to preserve the essential information of the hair effectively. This underscores the significance of incorporating both the scalp style details and the hair content information in the scalp disorder classification.
4 Conclusion and Discussion
In this work, we introduced ScalpVision, a diagnostic system designed for a complete evaluation of scalp health. Our approach combines label-free hair segmentation – based on a naive segmentation model and a foundation segmentation model (SAM) – with diffusion-based data augmentation to address data imbalance and preserve critical hair features. However, scalp disorders are affected by both the condition of the scalp and hair characteristics. Therefore, we plan to incorporate hair information to broaden our research to conditions such as alopecia, beyond the three scalp diseases. We see ScalpVision as an important step toward a more general diagnostic system for dermatological applications.
{credits}
4.0.1 Acknowledgements
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2024-00354218 and RS-2024-00353125). Junhyug Noh was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2022-00155966).
Appendix 0.A Detailed Overview of Scalp Diseases
[…]
Dandruff, also referred to as a milder manifestation of seborrheic dermatitis, is characterized by the non-inflammatory exfoliation of dead epidermal cells from the scalp. While it can induce mild itching, it generally does not precipitate erythema or the formation of scabs [3].
Hyperseborrhea, the excessive production of sebum, represents a common aesthetic concern, manifested through the secretion of excess oil from hypertrophic sebaceous glands. This condition results in a shiny, oily skin appearance. Although sebum plays a crucial role in maintaining skin hydration and its protective barrier, its excessive secretion can lead to various dermatological issues. One such issue is the formation of sebum plugs, which are small, yellowish, or pale bumps that appear on the skin [26].
Scalp erythema, also known as red scalp, is characterized by widespread redness across the scalp. It can arise from several conditions, including psoriasis, seborrheic dermatitis, contact dermatitis, diffuse lichen planopilaris, dermatomyositis, and scalp rosacea [29].
Appendix 0.B Implementation Details
Hair segmentation. For our heuristic-driven hair segmentation, we utilized the U2-Net [25] official source code to train 3,0003{,}000 pseudo image-label pairs. The training parameters included a batch size of $32$, a constant learning rate of $0.001$, $100$ training epochs, and the Adam optimizer. The output from U2-Net was binarized to obtain $\hat{M}$, using a threshold of $0.5$.
Since other previous studies, excluding SAM [15], do not have an official codebase, we implemented their approaches based on descriptions in their papers, using OpenCV. For instance, [28] used contrast stretching and binary thresholding to derive the hair mask. [32] applied morphological operations, while [13] used Otsu’s method [22] for hair mask acquisition. In evaluating SAM, the mask with the highest Intersection over Union (IoU) score was selected as the final prediction.
Image augmentation. For DiffuseIT-M, we construct a loss function for image translation while preserving hair content. This function is defined as follows:
$\begin{gathered}\ell_{total}(x;x_{src},x_{trg},M)=\lambda_{1}\ell_{style}+\lambda_{2}\ell_{content}\\ +\lambda_{3}\ell_{mask}+\lambda_{4}\ell_{sem}+\lambda_{5}\ell_{rng}.\end{gathered}$ (6)
To incorporate the semantic information of the target image, we establish the style loss function, $\ell_{sty}$. This function leverages the [CLS] token from the last layer of DINO-ViT [4]. Denoting the [CLS] tokens as $\mathbf{c}$, the style loss is expressed as:
[…]
where $\lambda_{mse}$ is set to 3,0003{,}000, and the weight for $\ell_{style}$, $\lambda_{1}$, is set to 2,0002{,}000.
Figure 6: Examples of pseudo images and their corresponding masks for hair segmentation.
The content loss, $\ell_{content}$, is designed to preserve the structure of source images. Let $k^{l}_{i}(x)$ represent the $i$-th key extracted from the $l$-th multi-head self-attention layer in DINO-ViT for image $x$. The content loss is then defined as:
$\ell_{content}=\lambda_{sim}\ell_{sim}(x_{src},\hat{x}_{0}(x_{t}))+\lambda_{con}\ell_{con}(x_{src},\hat{x}_{0}(x_{t})),$ (8)
where the similarity loss, $\ell_{sim}$, and the content loss, $\ell_{con}$, are
$\ell_{sim}(x_{src},\hat{x}_{0}(x_{t}))=||\text{cos}_{ij}(x_{src}),\text{cos}_{ij}(\hat{x}_{0}(x_{t}))||_{2},$ (9)
$\ell_{con}(x_{src},\hat{x}_{0}(x_{t}))=\text{infoNCE}(k^{l}_{i}(x_{src}),k^{l}_{i}(\hat{x}_{0}(x_{t}))),$ (10)
with $cos_{ij}(x)$ representing the cosine distance between $k^{l}_{i}(x)$ and $k^{l}_{j}(x)$. The weights $\lambda_{sim}$ and $\lambda_{con}$ are set to 1,0001{,}000 and $200$, respectively. Additionally, weights $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are set to $1,000$, $100$, and $200$, respectively. During our experiment, the model is configured to generate images with a resolution of $256\times 256$ pixels, utilizing a denoising step of 1,0001{,}000.
For the implementation of previous studies, we used the official source code of DiffuseIT [16] and AGG [17], conducting experiments with their provided hyperparameters. Non-generative image augmentations were performed using AugMix [10] through the torchvision library and Gaussian noise augmentation via the PyTorch library.
Scalp disease and severity diagnosis. For the classification task, we fine-tuned two models: DenseNet169 [12] (CNN-based) and EfficientFormerV2 [18] (Transformer-based), using pre-trained weights from the timm library [30]. Fine-tuning involved a batch size of $128$, a learning rate of $0.0001$ with a CosineAnnealingWarmRestarts scheduler [20], $50$ training epochs, and the AdamW optimizer [21].
Appendix 0.C Pseudo Image and Mask Visualization
To create a diverse pseudo training set, we extracted scalp patch images from areas without hair in nine different scalp images. Each image represented a unique disease at a specific severity level. As illustrated in Figure 6, we introduced a variety of hair types by inserting straight and curved lines in blue, brown, black, and white colors, each with differing thicknesses. To simulate common scalp noise, such as dandruff, white circular elements were added to the pseudo images. Our codebase contains further details on this process.
Appendix 0.D Scalp Disease and Severity Classification
This section outlines our data augmentation approach for classifying scalp diseases and their severities and presents additional experimental results.
1 Input: severitiesseverities: A collection of records for each severity. Assume that each element of severitiesseverities is classified as 0 (good), 1 (mild), 2 (moderate), and 3 (severe).
2Output: ratiosratios: Sampling ratio among four severity levels.
31:  $\epsilon\leftarrow 1\times 10^{-9}$
2:  sevCountssevCounts $\leftarrow$ $\{\}$
3:  for each sevLevelsevLevel in $[0,1,2,3]$ do
4:   $sevCounts[sevLevel]$ $\leftarrow$ 0
45:  end for
6:  for each severityseverity in severitiesseverities do
57:   $sevCounts[severity]$ $\leftarrow$ $sevCounts[severity]+1$
8:  end for
69:  invCountsinvCounts $\leftarrow$ $\{\}$
10:  for each sevLevelsevLevel in $[0,1,2,3]$ do
11:   $invCounts[sevLevel]$ $\leftarrow$ $1/(sevCounts[sevLevel]+\epsilon)$
712:  end for
813:  normFactornormFactor $\leftarrow$ $\sum_{sevLevel=0}^{3}invCounts[sevLevel]$
14:  ratiosratios $\leftarrow$ $[]$
15:  for each sevLevelsevLevel in $[0,1,2,3]$ do
16:   $ratio\leftarrow invCounts[sevLevel]/normFactor$
17:   $ratios[sevLevel]$ $\leftarrow$ ratioratio
18:  end for
19:  return ratiosratios
Algorithm 2 Calculation of sampling ratios for each disease
Figure 7: Distribution of data augmented using DiffuseIT-M. For comparison, the original data distribution is presented in Figure 2.
0.D.1 Data Augmentation Strategy
Addressing the issue of data imbalance in the dataset, we implemented a strategy to translate randomly selected images into classes with fewer samples. We used random selection for the source images and weighted sampling for target images, where the likelihood of choosing an image was inversely proportional to the number of samples in its severity class. This method favored the selection of underrepresented classes. The algorithm to calculate these sampling weights is detailed in Algorithm 2.
Figure 7 displays the data distribution after this augmentation. The post-augmentation distribution shows a more balanced representation across various classes, especially a rise in the severe category and a reduction in the good category. This balanced dataset played a crucial role in enhancing our model’s performance by providing an even distribution of the training samples. The same augmentation strategy was applied across all diffusion-based methods.
Table 4: Quantitative comparison of classification performance across different backbone models. “Baseline” refers to performance without any augmentation methods, while results following the “+” symbol indicate the use of various augmentation methods. Values in the table represent the macro-F1 scores.
Method CNN Transformer
ResNet ResNeXt EfficientFormer ViT
Baseline $0.5709$ $0.6186$ $0.5300$ $0.4934$
+ Gaussian Noise $0.4304$ $0.5663$ $0.5742$ $0.4306$
+ AugMix $0.4186$ $0.5247$ $0.5348$ $0.4747$
+ DiffuseIT $0.6064$ $0.6107$ $0.6120$ $0.5767$
+ AGG $0.5978$ $0.6251$ $0.6104$ $0.5771$
+ Ours 0.6128 0.6292 0.6170 0.5861
Figure 8: Qualitative results of DiffuseIT-M. This figure illustrates the results for various scalp disease conditions, with severity levels indicated as 0 (good), 1 (mild), 2 (moderate), and 3 (severe). Scalp diseases are color-coded for clarity: blue represents dandruff, green signifies excess sebum, and red denotes erythema.
Impact of backbone model. We assessed the performance of scalp disease classification using different pretrained backbone models. Our evaluation metric was the F1 macro score, which we also used to compare our results with existing augmentation methods, including those utilizing DiffuseIT and AGG. The experiments involved two CNN-based models (ResNet [9] and ResNeXt [31]) and two Transformer-based models (ViT [7] and EfficientFormer [19]), all maintaining consistent hyperparameters as described in Section 6.
As presented in Table 4, our approach achieved superior classification performance across all backbone models. Moreover, models augmented using generative methods consistently outperformed those using baseline augmentation, validating the effectiveness of our proposed data augmentation strategy.
Appendix 0.E Qualitative Results of DIffuseIT-M
Figure 8 presents qualitative results of our model’s ability to translate images across multiple labels while maintaining hair information. The results indicate that our model successfully translates various disease features from target images, effectively preserving hair representation.